OpenStudy (anonymous):

What is the value of –6 • 4 + (–7) • 3?

OpenStudy (anonymous):

wait a second sry lol i got it

OpenStudy (anonymous):

first do the multiplication: -6*4=-24 -7*3=-21 now do the additon -24 + (-21) = -24 - 21 = -45